OpenStudy (ellaella02):

What is the slope of a line whose equation is -48=2x-8y

OpenStudy (ellaella02):

@mathway

OpenStudy (anonymous):

Re-arrange it first. Make the equation look like: y=mx+b.

OpenStudy (ellaella02):

8y=2x+-48

OpenStudy (anonymous):

Then divide 8 on both sides. :)

OpenStudy (anonymous):

8y=2x+48**

OpenStudy (ellaella02):

Okay so it would be y=2x+6

OpenStudy (anonymous):

Quite right, but you did something wrong in the slope.

OpenStudy (anonymous):

48/8= 6 ((check)) 2/8=????

OpenStudy (ellaella02):

2 1 -- = --- 8 4

OpenStudy (anonymous):

Correct! :D Great job!

OpenStudy (ellaella02):

Is that the answer? 1/4

OpenStudy (ellaella02):

THANK YOU SO MUCH IT WAS RIGHT!!!!!